OpenStudy (anonymous):

Can you solve the problem and write out your steps. Then, use your work to write a detailed explanation of the steps you took to solve the problem. Make sure you use proper terminology, please. 3(7t + -3) = -4t + 41

OpenStudy (anonymous):

expand rhs

ganeshie8 (ganeshie8):

Step 1 : 3(7t - 3) = -4t + 41 Step 2 : 3 (7t) - 3 ( 3) = -4t + 41 => 21t - 9 = -4t + 41 Step 3 : 21t - 9 + 4t = -4t + 41 + 4t => 25t - 9 = 41 Step 4 : 25t -9 + 9 = 41 +9 => 25t = 50 Step 5 : 25t/25 = 50/25 => t = 2

ganeshie8 (ganeshie8):

Step 1 : Given equation Step 2 : Using distributive property of subtraction, a(b-c) = ab-ac Step 3 : Adding 4t to both sides Step 4 : Adding 9 to both sides Step 5 : Dividing both sides by 25